# Rethinking Self-training for Semi-supervised Landmark Detection: A Selection-free Approach

### v1

@@ Abstract
Landmark detection is a vision task that aims to localize predefined points on a given image. Its application covers several areas such as facial [1, 2] and medical landmark detection [3, 4, 5], both of which can be the preceding steps to other tasks (e.g., face recognition [6], face editing [7], surgery planning [8, 9], etc.). Thus, it is essential for a landmark model to accurately and robustly localize points in the wild.
@@
Based on the above observation, we propose Self-Training for Landmark Detection (STLD), a method that does not rely on explicit pseudo-label selection. Instead, it leverages pseudo-labels by constructing a task curriculum, where confident tasks are conducted first for obtaining reliable information from the pseudo-labels. Specifically, the curriculum starts with pseudo pretraining by only leveraging pseudo-labels for providing a better initialization; then shrink regression is further added to the curriculum in the later rounds to directly leverage the pseudo-labels via a coarse-to-fine process. The process starts with coarse-grained regression, then progressively increases the granularity over the rounds until the same level as standard regression. In each round, STLD follows a two-stage pipeline. The first stage conducts pseudo pretraining on all the pseudo-labeled data; the second stage, termed source-aware mixed training, leverages ground-truths (GTs) and pseudo-labeled data via standard and shrink regression, respectively. The proposed method is model-agnostic and works with both heatmap and coordinate regression. Fig. 1 shows the training pipeline of STLD. Our experiments on three facial and one medical landmark detection benchmark demonstrate the superiority of the proposed method, showing that STLD surpasses the existing SSL methods in both semi- and omni-supervised settings. For example, in the semi-supervised setting, our method improves the performance over SOTA methods by 5.7% on WFLW; in the omni setting, STLD outperforms its counterparts with less unlabeled images (200K 20K) on 300W.
@@ II-C Semi-supervised Landmark Detection
Semi-supervised landmark detection can be divided into three types. 1) Consistency regularization based methods force models to learn consistency across different geometric [50, 51] or style [52] transformations to obtain supervision for unlabeled images. 2) Some methods attempt to learn more generalizable representations [53] or reasonable predictions [48] with unlabeled data through adversarial learning. 3) Self-training based methods iteratively estimate and utilize pseudo-labels of unlabeled data. To address confirmation bias in self-training [54], Dong et al. [21] designed a teacher network to assess the quality of pseudo-labels for sample selection, but it only works with small input size (64×64) due to training difficulty. Jin et al. [55] proposed a self-training method to gradually increase the difficulty of a simpler proxy task, which does not require pseudo-label selection. However, their method [55] is closely tied to a specific landmark detector, thus it is inapplicable to other SOTA models. HybridMatch [56] proposed to combine 1-D and 2-D heatmap regression to alleviate the task-oriented and training-oriented issues, respectively, based on FixMatch [28]. Compared to these existing works [21, 55, 56], we explicitly tackle the incompatibility of self-training for landmark detection by proposing a generic method that does not rely on explicit pseudo-label filtering.
@@ III-B Selection-based Self-training
Although selecting confident pseudo-labels mitigates the confirmation bias, it may introduce data bias. To verify this, we analyzed the label distribution of different data groups via label density map (see Fig. 2b for visualized label density maps). By calculating the KL divergence of the label density maps of two data groups, we could quantify the discrepancy of the two groups. The analysis was based on 300W [63], where we used the GT of the unlabeled data as the anchor to calculate its KL divergence with three data groups respectively: 1) the labeled data (1.6%); 2) all the unlabeled data with pseudo-labels estimated by the model trained on the labeled data; 3) unlabeled data with confident pseudo-labels (). Fig. 2a shows the KL divergence between the unlabeled GT and the three groups averaged over all the landmarks as well as the ones of several randomly selected landmarks. It can be seen from the figure that the labeled data has the largest KL divergence to the unlabeled GT on average, indicating that limited number of labeled data can be quite biased and thus limits the generalization of the model. The confident pseudo-labels have a smaller KL divergence on average compared to the labeled data (1.69 vs. 2.70), yet its distribution shift is still significant. Interestingly, the average KL divergence of all the pseudo-labels is much smaller than that of the other two groups (0.45 vs. 1.69/2.70). In other words, selecting confident pseudo-labels reduces the errors from noisy pseudo-labels, but also introduces data bias due to the biased selection. Additionally, it is not easy to decide a proper threshold for sample selection, especially for the fine-grained localization task that is more sensitive to noisy labels. Moreover, coordinate-regression based methods usually output predictions without confidence scores, which makes classic self-training inapplicable. Thus, we pose the natural question: Is pseudo-label selection a necessary component of self-training?
@@ IV-A Pseudo Pretraining for Better Initialization
During the pseudo pretraining stage, i.e., the simplest task of the curriculum, model is not asked to fit the pseudo-labels directly, but rather obtaining a better initialization (compared to random) with them. To this end, the model is simply pretrained on all the pseudo-labeled data with the following formula:
@@ IV-B Shrink Regression for Noise Resistance
Although pseudo pretraining handles noisy samples well, it has not fully leveraged the pseudo-labels as they are only used for model initialization. To achieve better performance, it is necessary to utilize pseudo-labels directly for model training while taking care of the noise. To this end, we first inspect the characteristics of noisy labels by visualizing the offsets of pseudo-labels relative to the GTs in 2D histograms, as shown in Fig. 4. The figure indicates that pseudo-labels are distributed around GTs for both heatmap (Figs. 4a and 4b) and coordinate regression (Figs. 4c and 4d) across different labeled ratios. For those pseudo-labels that are relatively far from the GTs, if the model directly fits to them with the standard granularity, it would learn the noisy information introduced by these pseudo-labels due to their deviation from GTs. On the other hand, if a coarse granularity is used for regression, the model fits to a broader region centered on the pseudo-labels, which is more likely to cover the GTs. Accordingly, we argue that the standard regression task on pseudo-labels can be made more confident by reducing the granularity of regression targets. In other words, it sacrifies the precision of regression for the reliability of learning targets. Following this observation, we propose shrink regression, which starts with coarse regression, then progressively increases the granularity until the same level as standard regression. When it is applied, the in Eqn. 4 can be implemented as the loss function , where represents the -th round and . Here starts from 2 because shrink regression is applied from the second round. Different instantiations of for heatmap and coordinate regression are detailed in Sec. IV-B1 and Sec. IV-B2, respectively. A supervised landmark detection model can be seen as a special case of shrink regression, where its granularity is irrelevant to and fixed to the standard value.
@@ IV-C Training Pipeline
| Labeled ratio |
1.6% |
5% |
10% |
20% |
0.7% |
5% |
10% |
20% |
1.0% |
5% |
10% |
20% |
| RCN+ [50] |
5.11 |
4.47 |
4.14 |
2.88 |
2.17 |
| TS [21] |
5.64 |
5.03 |
2.19 |
2.14 |
1.99 |
| 3FabRec [53] |
5.10 |
4.75 |
4.47 |
4.31 |
8.39 |
7.68 |
6.73 |
6.51 |
2.38 |
2.13 |
2.03 |
1.96 |
| STC [55] |
4.72 |
4.04 |
3.76 |
3.58 |
7.99 |
5.94 |
5.40 |
5.09 |
2.05 |
1.79 |
1.74 |
1.68 |
| PL [15] |
5.45 |
4.33 |
3.80 |
3.60 |
9.04 |
6.00 |
5.47 |
5.14 |
2.20 |
1.85 |
1.78 |
1.70 |
| Data Dist. [32] |
5.38 |
4.35 |
3.82 |
3.59 |
8.99 |
6.06 |
5.50 |
5.18 |
2.18 |
1.86 |
1.80 |
1.72 |
| CL [19] |
5.61 |
4.16 |
3.76 |
3.58 |
8.87 |
5.98 |
5.45 |
5.14 |
2.13 |
1.83 |
1.75 |
1.70 |
| Easy-hard Aug. [51] |
5.00 |
4.18 |
3.78 |
3.60 |
8.63 |
5.90 |
5.41 |
5.14 |
2.10 |
1.80 |
1.75 |
1.69 |
| FixMatch [28] |
4.94 |
4.15 |
3.76 |
3.59 |
8.57 |
5.86 |
5.39 |
5.12 |
2.09 |
1.79 |
1.74 |
1.69 |
| DST [66] |
4.66 |
4.02 |
3.75 |
3.57 |
7.95 |
5.81 |
5.36 |
5.10 |
2.06 |
1.78 |
1.73 |
1.67 |
| HybridMatch [56] | 3.76 | 3.50 |
3.44 |
3.40 |
5.47 | 5.04 |
4.85 |
4.73 |
1.77 | 1.68 |
1.66 |
1.61 |
| STLD-HM-R18 (Ours) |
4.47 |
3.81 |
3.64 |
3.52 |
7.91 |
5.67 |
5.30 |
5.06 |
2.03 |
1.75 |
1.69 |
1.63 |
| STLD-TF-R18 (Ours) |
4.97 |
4.00 |
3.69 |
3.49 |
8.07 |
5.68 |
5.17 |
4.82 |
2.10 |
1.76 |
1.67 |
1.60 |
@@
| STLD-TF-HRNet (Ours) |
4.35 |
3.69 |
3.42 | 3.25 |
6.65 |
5.06 |
4.77 | 4.46 |
1.96 |
1.65 | 1.60 | 1.53 |
@@ V-A2 Medical landmark
The Hand is a public dataset111https://ipilab.usc.edu/research/baaweb of hand X-ray images. There are 609 and 300 images for training and testing, respectively. We resize each image to for preprocessing. Payer et al. [71] labeled 37 landmarks for each image and assume 50mm wrist width as physical distance, which is adopted in this paper. Additionally, we use 12K images from RSNA [72] training set as the extra unlabeled data for omni-supervised learning.
@@ V-B1 Implementation details
To demonstrate the universality of our method, we use two base models, one is heatmap-based and the other is coordinate-based. The heatmap model (HM) is adopted from [64] and the coordinate model uses transformer decoder as the head (TF), adopted from [65]. We use ImageNet pretrained ResNet-18 [73] as the default backbone, unless otherwise stated. The initial learning rate is 1e-4, and Adam [74] is used as the optimizer. The batch size is set to 16. At each training stage, the total training epoch is 60 for HM, learning rate decayed by 10 at 40th and 50th epoch; the training epoch is 360 for TF, decayed by 10 at 240th and 320th epoch. For standard regression, HM uses and TF uses (i.e., loss), following common practice [47, 65]. The code was implemented with PyTorch 1.13 and trained with one RTX 3090 GPU. The training of STLD on 300W with 5% labeled takes about 3 and 12 hours for HM and TF, respectively.
@@ V-B4 Evaluation metric
AUC |
FR |
AUC |
FR |
| CL [19] |
62.26 |
0.73 |
64.46 |
0.44 |
| FixMatch [28] |
62.34 |
0.87 |
64.06 |
0.87 |
| DST [66] |
62.58 |
0.44 |
64.13 |
0.29 |
| HybridMatch [56] |
60.56 |
0.17 |
| STLD-HM-R18 (Ours) |
63.30 |
0.44 |
65.26 |
0.29 |
| STLD-TF-R18 (Ours) |
62.79 |
0.58 |
65.26 |
0.29 |
@@
| STLD-TF-HRNet (Ours) |
65.64 |
0.29 |
67.44 |
0.15 |
| Method | Backbone |
Unlabeled |
NME (%) |
| TS [21] | HG+CPM |
20K |
3.49 |
| LaplaceKL [48] | - |
70K |
3.91 |
| DeCaFA [75] | U-net |
200K |
3.39 |
| STC [55] | ResNet-18 |
200K |
3.23 |
| Data Dist. [32] | ResNet-18 |
200K |
3.31 |
| CL [19] | ResNet-18 |
200K |
3.26 |
| Easy-hard Aug. [51] | ResNst-18 |
200K |
3.27 |
| STLD-HM (Ours) | ResNet-18 |
20K |
3.29 |
70K |
3.27 |
200K |
3.23 |
| STLD-TF (Ours) | ResNet-18 |
20K |
3.20 |
70K |
3.18 | ||
200K |
3.14 |

### v2

@@ Abstract
Landmark detection is a vision task that aims to localize predefined points on a given image. Its application covers several areas such as facial [1, 2, 3] and medical landmark detection [4, 5, 6], both of which can be the preceding steps to other tasks (e.g., face recognition [7], face editing [8], surgery planning [9, 10], etc.). Thus, it is essential for a landmark model to accurately and robustly localize points in the wild.
@@
Based on the above observation, we propose Self-Training for Landmark Detection (STLD), a method that does not rely on explicit pseudo-label selection. Instead, it leverages pseudo-labels by constructing a task curriculum, where confident tasks are conducted first for obtaining reliable information from the pseudo-labels. Specifically, the curriculum starts with pseudo pretraining by only leveraging pseudo-labels for providing a better initialization; then shrink regression is further added to the curriculum in the later rounds to directly leverage the pseudo-labels via a coarse-to-fine process. The process starts with coarse-grained regression, then progressively increases the granularity over the rounds until the same level as standard regression. In each round, STLD follows a two-stage pipeline. The first stage conducts pseudo pretraining on all the pseudo-labeled data; the second stage, termed source-aware mixed training, leverages ground-truths (GTs) and pseudo-labeled data via standard and shrink regression, respectively. The proposed method is model-agnostic and works with both heatmap and coordinate regression. Fig. 1 shows the training pipeline of STLD. Our experiments on three facial and one medical landmark detection benchmark demonstrate the superiority of the proposed method, showing that STLD surpasses the existing SSL methods in both semi- and omni-supervised settings. For example, in the semi-supervised setting, our method improves the performance over SOTA methods by 5.7% on WFLW; in the omni setting, STLD outperforms its counterparts with less unlabeled images (200K 20K) on 300W. We summarize our contributions as follows.
We propose to rethink self-training for semi-supervised landmark detection from a different perspective by changing from the selection-based data-level curriculum to selection-free task-level curriculum. The proposed task-level curriculum has the potential to address the problems of previous methods, namely 1) the data bias issue due to sample selection, 2) the sensitivity to threshold, and 3) the infeasibility to coordinate regression-based methods.
Following the philosophy of task-level curriculum, we design a novel framework for semi-supervised landmark detection. The framework gradually increases the granularity of the task over self-training rounds, enabling the model to leverage the knowledge of pseudo-labels while introducing minimum noise. Pseudo pretraining is designed as the task with the lowest level of granularity by pretraining the model on all the pseudo-labeled data; shrink regression is further added in the later rounds to directly leverage the pseudo-labels by conducting the regression task in a coarse-to-fine manner. The proposed framework is model-agnostic and works with both heatmap and coordinate regression.
Extensive experiments are conducted to demonstrate the effectiveness of the proposed framework in both semi- and omni-supervised settings, where the datasets cover one medical and three facial landmark detection benchmarks.
@@ II-C Semi-supervised Landmark Detection
Semi-supervised landmark detection can be divided into three types. 1) Consistency regularization based methods force models to learn consistency across different geometric [51, 52] or style [53] transformations to obtain supervision for unlabeled images. 2) Some methods attempt to learn more generalizable representations [54] or reasonable predictions [49] with unlabeled data through adversarial learning. 3) Self-training based methods iteratively estimate and utilize pseudo-labels of unlabeled data. To address confirmation bias in self-training [55], Dong et al. [22] designed a teacher network to assess the quality of pseudo-labels for sample selection, but it only works with small input size (64×64) due to training difficulty. Jin et al. [56] proposed a self-training method to gradually increase the difficulty of a simpler proxy task, which does not require pseudo-label selection. However, their method [56] is closely tied to a specific landmark detector, thus it is inapplicable to other SOTA models. HybridMatch [57] proposed to combine 1D and 2D heatmap regression to alleviate the task-oriented and training-oriented issues, respectively, based on FixMatch [29]. Additionally, a confidence-based soft filtering strategy is proposed in HybridMatch to avoid thresholding by regularizing the pseudo-labeled loss with prediction confidence, which makes it also a selection-free method. However, due to the dependence on prediction confidence, HybridMatch would inevitably introduce noise as modern neural networks are rarely calibrated and tend to give overconfident predictions [58, 59]. Compared to these existing works [22, 56, 57], our method is a more general solution that 1) applies to both heatmap and coordinate regression, and 2) relies on neither sample selection nor confidence regularization. We attribute these advantages to the carefully designed task-level curriculum.
@@ III-B Selection-based Self-training
Although selecting confident pseudo-labels mitigates the confirmation bias, it may introduce data bias. To verify this, we analyze the label distribution of different data groups via label density map. The analysis is based on 300W [66]. There are four groups of labels: 1) the GT of unlabeled data (98.4%); 2) the labeled data (1.6%); 3) all the unlabeled data with pseudo-labels estimated by the model trained on the labeled data; 4) unlabeled data with confident pseudo-labels (). To obtain label density maps, we first divide the 256256 map into 1212 bins, where the downsampling from 256 to 12 is to avoid the impact of small changes so that it is shift invariant to a certain degree. Then the label density maps of different groups can be obtained by calculating the frequency of landmarks within each bin, followed by normalization. Fig. 2a visualizes the label density maps of two landmarks with ID 64 and 56. It could be visually perceived that the density map of all pseudo is more similar to that of unlabeled GT than the other two groups. To quantify the similarity, we use KL divergence to measure the distance between the anchor distribution (i.e., unlabeled GT) and the distribution of the remaining groups, respectively. KL divergence is used because it is a standard measure of the dissimilarity between two probability distributions [67].
Fig. 2b shows the KL divergence between the unlabeled GT and the other three groups averaged over all the landmarks as well as the ones of several randomly selected landmarks. It can be seen from the figure that the labeled data has the largest KL divergence to the unlabeled GT on average, indicating that limited number of labeled data can be quite biased and thus limits the generalization of the model. The confident pseudo-labels have a smaller KL divergence on average compared to the labeled data (1.69 vs. 2.70), yet its distribution shift is still significant. Interestingly, the average KL divergence of all the pseudo-labels is much smaller than that of the other two groups (0.45 vs. 1.69/2.70). In other words, selecting confident pseudo-labels reduces the errors from noisy pseudo-labels, but also introduces data bias due to the biased selection. Additionally, it is not easy to decide a proper threshold for sample selection, especially for the fine-grained localization task that is more sensitive to noisy labels. Moreover, coordinate-regression based methods usually output predictions without confidence scores, which makes classic self-training inapplicable. Thus, we pose the natural question: Is pseudo-label selection a necessary component of self-training?
@@ IV-A Pseudo Pretraining for Better Initialization
During the pseudo pretraining stage, i.e., the simplest task of the curriculum, a model is not asked to fit the pseudo-labels directly, but rather obtaining a better initialization with them. To this end, the model initialized from ImageNet [68] pretrained weights is trained on all the pseudo-labeled data with the following formula:
@@ IV-B Shrink Regression for Noise Resistance
Although pseudo pretraining handles noisy samples well, it has not fully leveraged the pseudo-labels as they are only used for model initialization. To achieve better performance, it is necessary to utilize pseudo-labels directly for model training while taking care of the noise. To this end, we first inspect the characteristics of noisy labels by visualizing the offsets of pseudo-labels relative to the GTs in 2D histograms, as shown in Fig. 4. The figure indicates that pseudo-labels are distributed around GTs for both heatmap (Figs. 4a and 4b) and coordinate regression (Figs. 4c and 4d) across different labeled ratios. For those pseudo-labels that are relatively far from the GTs, if the model directly fits to them with the standard granularity111The granularity here refers to the precision of regression task. A fine-grained regression task (i.e., high-level granularity) penalizes more on small prediction errors while a coarse-grained regression (i.e., low-level granularity) is more tolerant to small errors., it would learn the noisy information introduced by these pseudo-labels due to their deviation from GTs. On the other hand, if a coarse granularity is used for regression, the model fits to a broader region centered on the pseudo-labels, which is more likely to cover the GTs. Accordingly, we argue that the standard regression task on pseudo-labels can be made more confident by reducing the granularity of regression targets. In other words, it sacrifies the precision of regression for the reliability of learning targets. Following this observation, we propose shrink regression, which starts with coarse regression, then progressively increases the granularity until the same level as standard regression. When it is applied, the in Eqn. 4 can be implemented as the loss function , where represents the -th round and . Here starts from 2 because shrink regression is applied from the second round. Different instantiations of for heatmap and coordinate regression are detailed in Sec. IV-B1 and Sec. IV-B2, respectively. A supervised landmark detection model can be seen as a special case of shrink regression, where its granularity is irrelevant to and fixed to the standard value.
@@ IV-C Training Pipeline
| Labeled ratio | 1.6% | 5% | 10% | 20% | 0.7% | 5% | 10% | 20% | 1.0% | 5% | 10% | 20% |
| RCN+ [51] | - | 5.11 | 4.47 | 4.14 | - | - | - | - | 2.88 | 2.17 | - | - |
| TS3 [22] | - | - | 5.64 | 5.03 | - | - | - | - | - | 2.19 | 2.14 | 1.99 |
| 3FabRec∗ [54] | 5.10 | 4.75 | 4.47 | 4.31 | 8.39 | 7.68 | 6.73 | 6.51 | 2.38 | 2.13 | 2.03 | 1.96 |
| STC [56] | 4.72 | 4.04 | 3.76 | 3.58 | 7.99 | 5.94 | 5.40 | 5.09 | 2.05 | 1.79 | 1.74 | 1.68 |
| PL† [16] | 5.45 | 4.33 | 3.80 | 3.60 | 9.04 | 6.00 | 5.47 | 5.14 | 2.20 | 1.85 | 1.78 | 1.70 |
| Data Dist.† [33] | 5.38 | 4.35 | 3.82 | 3.59 | 8.99 | 6.06 | 5.50 | 5.18 | 2.18 | 1.86 | 1.80 | 1.72 |
| CL† [20] | 5.61 | 4.16 | 3.76 | 3.58 | 8.87 | 5.98 | 5.45 | 5.14 | 2.13 | 1.83 | 1.75 | 1.70 |
| Easy-hard Aug.† [52] | 5.00 | 4.18 | 3.78 | 3.60 | 8.63 | 5.90 | 5.41 | 5.14 | 2.10 | 1.80 | 1.75 | 1.69 |
| FixMatch† [29] | 4.94 | 4.15 | 3.76 | 3.59 | 8.57 | 5.86 | 5.39 | 5.12 | 2.09 | 1.79 | 1.74 | 1.69 |
| DST† [71] | 4.66 | 4.02 | 3.75 | 3.57 | 7.95 | 5.81 | 5.36 | 5.10 | 2.06 | 1.78 | 1.73 | 1.67 |
| HybridMatch [57] | 3.76 | 3.50 | 3.44 | 3.40 | 5.47 | 5.04 | 4.85 | 4.73 | 1.77 | 1.68 | 1.66 | 1.61 |
| STLD-HM-R18 (Ours) | 4.47 | 3.81 | 3.64 | 3.52 | 7.91 | 5.67 | 5.30 | 5.06 | 2.03 | 1.75 | 1.69 | 1.63 |
| STLD-TF-R18 (Ours) | 4.97 | 4.00 | 3.69 | 3.49 | 8.07 | 5.68 | 5.17 | 4.82 | 2.10 | 1.76 | 1.67 | 1.60 |
@@
| STLD-TF-HRNet (Ours) | 4.35 | 3.69 | 3.42 | 3.25 | 6.65 | 5.06 | 4.77 | 4.46 | 1.96 | 1.65 | 1.60 | 1.53 |
@@ V-A2 Medical landmark
The Hand is a public dataset222https://ipilab.usc.edu/research/baaweb of hand X-ray images. There are 609 and 300 images for training and testing, respectively. We resize each image to for preprocessing. Payer et al. [76] labeled 37 landmarks for each image and assume 50mm wrist width as physical distance, which is adopted in this paper. Additionally, we use 12K images from RSNA [77] training set as the extra unlabeled data for omni-supervised learning.
@@ V-B1 Implementation details
To demonstrate the universality of our method, we use two base models, one is heatmap-based and the other is coordinate-based. The heatmap model (HM) is adopted from [69] and the coordinate model uses transformer decoder as the head (TF), adopted from [70]. ResNet-18 [78] is used as the default backbone and HRNet [48] is also used to explore better performance. Both networks are initialized from ImageNet [68] pretrained weights. The initial learning rate is 1e-4, and Adam [79] is used as the optimizer. The batch size is set to 16. At each training stage, the total training epoch is 60 for HM, learning rate decayed by 10 at 40th and 50th epoch; the training epoch is 360 for TF, decayed by 10 at 240th and 320th epoch. For standard regression, HM uses and TF uses (i.e., loss), following common practice [48, 70]. The code was implemented with PyTorch 1.13.
@@ V-B4 Evaluation metric
| AUC | FR | AUC | FR | |
| CL† [20] | 62.26 | 0.73 | 64.46 | 0.44 |
| FixMatch† [29] | 62.34 | 0.87 | 64.06 | 0.87 |
| DST† [71] | 62.58 | 0.44 | 64.13 | 0.29 |
| HybridMatch [57] | - | - | 60.56 | 0.17 |
| STLD-HM-R18 (Ours) | 63.30 | 0.44 | 65.26 | 0.29 |
| STLD-TF-R18 (Ours) | 62.79 | 0.58 | 65.26 | 0.29 |
@@
| STLD-TF-HRNet (Ours) | 65.64 | 0.29 | 67.44 | 0.15 |
| Method | Backbone | Unlabeled | NME (%) |
| TS3 [22] | HG+CPM | 20K | 3.49 |
| LaplaceKL [49] | - | 70K | 3.91 |
| DeCaFA [80] | U-net | 200K | 3.39 |
| STC [56] | ResNet-18 | 200K | 3.23 |
| Data Dist.† [33] | ResNet-18 | 200K | 3.31 |
| CL† [20] | ResNet-18 | 200K | 3.26 |
| Easy-hard Aug.† [52] | ResNst-18 | 200K | 3.27 |
| STLD-HM (Ours) | ResNet-18 | 20K | 3.29 |
| 70K | 3.27 | ||
| 200K | 3.23 | ||
| STLD-TF (Ours) | ResNet-18 | 20K | 3.20 |
| 70K | 3.18 | ||
| 200K | 3.14 |